# Rethinking Weight-Averaged Model-merging

### v1

@@ 6.3 Model Predictions and Weight Magnitudes
Specifically, we visualize the merging of 10 models (merging of 2 to 7 models can be found in the supplementary materials) for model-merging and ensemble towards different magnitude factors, as shown in Fig. 8(b). In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with weight magnification. From the trends in the figures, we can observe the model-merging models are more robust than model ens models, including ensemble on logits and features. But surprisingly, feature ensemble are more robust than logits ensemble across different configurations. Feature ensemble models can even outperform model-merging when weight magnification is not large (e.g., ×$\times 50$). This shows the feature ensemble can have some resistance towards the change of weight magnitudes and variances. The “Ens Features” model uses the merge of last fully connected (FC) layers from all merged models for classification. In order to mitigate the impact of the merged FC layer, we introduce a variant of the model, “Ens Features*”, by directly adopting the FC layer from the 1st model. We can observe the “Ens Features” and “Ens Features*” models perform similarly.
We further visualize the gaps between weight-averaged model-merging and ensemble in Fig. 8(c). When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ for Cifar100). The same experiments to visualize the merging of 2 to 10 models for Cifar10 with different magnitude factors, as well as the gaps between weight-averaged model-merging and ensemble, are shown in the supplementary materials. Also, a similar phenomenon is shown on Cifar10, but with different peaks of gaps w.r.t. the magnitudes. Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab. 3. From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent. It can also outperform logit ensemble, but can not improve over ensemble on features and is not as good as on Cifar100 (e.g. 4.81% to 6.86% accuracy for 10 models merging). This may be caused the complexity of the Tiny ImageNet task and as we discussed in the Sec. 5.3, the model-merging can work to some extent as it performs similar operations as regularization (e.g., L1 or L2 regularization). It will mitigate the weight magnitude changing issues, but with the cost of losing model expressiveness.
Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$. This approach works for smaller data (Cifar10/Cifar100) as it plays aregularization role. However it does not work forlarger datasets as Tiny ImageNet, because it sacrifices the expressiveness of the models.
@@ 7 Conclusion
In the paper, we posit weight-averaged model-merging from a template matching perspective by visualizing weight patterns across different datasets. Additionally, our comparison between weight and feature averaging in ensemble strategies revealed distinctions in their impacts across diverse model architectures and datasets, providing valuable insights into when and why each approach can be advantageous. Finally, we found model-merge techniques can provide a prediction that is more stable than model ensemble techniques, particularly under different parameter magnitudes/variances, contributing to increased robustness and generalization. These insights clarify interpretability and critical aspects of model-merging for understanding and applying this technique in various contexts. We hope these findings will encourage further exploration into model-merging methods and inform future development of interpretable and robust model-averaging techniques.
@@ 8 Experimental Settings
For the experimental settings of model weight patterns. We train one single linear layer on both Cifar100 [12] and Tiny ImageNet [13] datasets with momentum SGD as the optimizer. The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1. While visualization, the weights are wrapped back to the image size (e.g. 100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [33] for weight-averaged model-merging. We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances. For the models considered, we have ResNet [8] (ResNet18 to ResNet152), DenseNet [9] (DenseNet121 to DenseNet201), VGG [22] (VGG11 to VGG19), ViT [24, 6] and DeiT [27]. For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [37, 38], Celeba [15], Tiny ImageNet and ChestXRay [31, 16]. The information of these datasets can be find in Tab. 4. Specially, ChestXRay dataset has 14 labels and each of them is a binary classification task. We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay. For model ensemble, we consider ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer). Besides these two model ensemble techniques and whenever a model is considered to be merged into a greedy soup model, we ensemble it using either the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts). In order to exclude the impact from adopting the last layer from any single model, the “Ens Features” model uses the merge of last fully connected (FC) layers from all merged models for classification.
@@
For all experiments in this section, to keep a fair comparison, we adopt the official splits for each dataset. Also, we uniformly train 10 models and perform the model-merging and ensemble under exactly same conditions.
@@ 9 More Visualization of Model Weight Patterns
Figure 9: Average of all images within each class (all 100 classes) of Cifar100 dataset.
Figure 10: Linear classifiers visualization for each class (all 100 classes) of Cifar100 dataset.
Similar to the visualization in the main paper, we visualize the average of all images Fig. 9 and linear classifiers for all classes Fig. 10 within each class (all 100 classes) of Cifar100 dataset. We can clearly observe the positive correlation between the average of images in a class and the learned weights.
@@
More visualization of the convolutional kernels on Cifar100 within deep learning models, e.g. Fig. 12 and Fig. 13.
Figure 14: Average of all images within each class (the first 100 classes) of TinyImageNet dataset.
Figure 15: Linear classifiers visualization for each class (the first 100 classes) of TinyImageNet dataset.
Figure 16: Average of all images within each class (the last 100 classes) of TinyImageNet dataset.
Figure 17: Linear classifiers visualization for each class (the last 100 classes) of TinyImageNet dataset.
@@
Figure 18: The 2 model merging/ensemble accuracy comparison between different architectures on Cifar100.
Figure 19: The 3 model merging/ensemble accuracy comparison between different architectures on Cifar100.
Figure 20: The 5 model merging/ensemble accuracy comparison between different architectures on Cifar100.
Figure 21: The 7 model merging/ensemble accuracy comparison between different architectures on Cifar100.
The 2 models, 3 models, 5 models and 7 models for model merging/ensemble accuracy comparison between different architectures on Cifar100 are shown via Fig. 18, Fig. 19, Fig. 20 and Fig. 21, respectively. They all show that the ensemble on logits generally receives the best results and uniform soups receives the worst results. Besides, model-merging performs much worse results for the ViT models and the gaps between uniform soups and ensemble on logits are enlarged with more models considered in.
Figure 22: Performance Gaps between Logits Ensemble and Greedy Soups Across Different Configurations (excluded the ViT models as the scales are too different) on Cifar100. It shows the greedy soups can mitigate the marginal effect problem to some extent.
From Fig. 22, we can observe similar with the gaps between ensemble on logits with uniform soups, with more models merged, the larger gaps between ensemble on logits with greedy soups will be. But differently, the greedy soups break the going down trend with uniform soups.
Table 5: Performance Comparison for VGG19 with 2 Models Across Different Datasets. We adopt averged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
@@
Table 6: Performance Comparison for VGG19 with 3 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
@@
Table 7: Performance Comparison for VGG19 with 5 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
@@
Table 8: Performance Comparison for VGG19 with 7 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay dataset; while accuracy (0% to 100%) for the rest of datasets.
@@
Table 9: Comparison of FLOPs and Parameters for Various Models. The FLOPs are calculated based on 224 $\times$ 224 for general purpose, e.g. ImageNet. “ViT_Tiny” denotes for “ViT_Tiny_Patch16_224” and “DeiT_Tiny” stands for “DeiT_Tiny_Patch16_224”, respectively.
@@
We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in Tab. 10. We can also see the more models to be merged, the poorer results from uniform soups can get. Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features bearly changed, it further demonstrates the model-merging between 2 or more models will not contribute to the performance.
@@ 11 More Experimental Results of Model Magnitude Changes
Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs. magnification factor on Cifar10 are shown via Fig. 27, Fig. 28, Fig. 29, Fig. 30 and Fig. 31, respectively. Moreover, we also visualize the uniform soup and logits ensemble accuracy vs. magnification factor across different model numbers on the Cifar10 dataset in Fig. 32. We can observe similar phenomenons. However, on the Cifar10, the gaps have different peak point, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).

### v2

@@ 6.3 Model Predictions and Weight Magnitudes
Specifically, we visualize the merging of 10 models (merging of 2 to 7 models can be found in the supplementary materials) for model-merging and ensemble towards different magnitude factors, as shown in Fig. 8(b). In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with weight magnification. From the trends in the figures, we can observe the model-merging models are more robust than model ens models, including ensemble on logits and features. But surprisingly, feature ensembles are more robust than logits ensembles across different configurations. Feature ensemble models can even outperform model merging when weight magnification is not large (e.g., ×$\times 50$). This shows the feature ensemble can have some resistance towards the change of weight magnitudes and variances. The “Ens Features” model uses the merge of the last fully connected (FC) layers from all merged models for classification. In order to mitigate the impact of the merged FC layer, we introduce a variant of the model, “Ens Features*”, by directly adopting the FC layer from the 1st model. We can observe the “Ens Features” and “Ens Features*” models perform similarly.
We further visualize the gaps between weight-averaged model-merging and ensemble in Fig. 8(c). When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ for Cifar100). The same experiments to visualize the merging of 2 to 10 models for Cifar10 with different magnitude factors, as well as the gaps between weight-averaged model-merging and ensemble, are shown in the supplementary materials. Also, a similar phenomenon is shown on Cifar10, but with different peaks of gaps w.r.t. the magnitudes. Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab. 3. From the table, we can observe that model merging can still mitigate the weight magnifying problem to a certain extent. It can also outperform logit ensembles, but can not improve over ensemble on features and is not as good as on Cifar100 (e.g. 4.81% to 6.86% accuracy for 10 models merging). This may be caused by the complexity of the Tiny ImageNet task and as we discussed in the Sec. 5.3, the model-merging can work to some extent as it performs similar operations as regularization (e.g., L1 or L2 regularization). It will mitigate the weight magnitude changing issues, but with the cost of losing model expressiveness.
Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$. This approach works for smaller data (Cifar10/Cifar100) as it plays a regularization role. However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
@@ 7 Conclusion
In this paper, we thoroughly examined the weight-averaged model-merging approach through the lens of template matching by visualizing weight patterns across various datasets. Our comparison of weight and feature averaging within ensemble and merging methods revealed notable differences in their effects on different model architectures and datasets, offering insights into when and why each approach is beneficial. Furthermore, we demonstrate that model merging techniques tend to yield more stable predictions than model ensembles, especially when subjected to variations in parameter magnitudes, thus enhancing robustness and generalization. These findings clarify the interpretability and key aspects of model-merging, providing a deeper understanding of its application across diverse contexts. We anticipate that these insights will stimulate further research into model merging strategies and support the development of more interpretable and robust model-averaging methodologies.
@@ 8 Experimental Settings
For the experimental settings of model weight patterns. We train one single linear layer on both Cifar100 [12] and Tiny ImageNet [13] datasets with momentum SGD as the optimizer. The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1. While visualization, the weights are wrapped back to the image size.
In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [33] for weight-averaged model-merging. We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances. For the models considered, we have ResNet [8] (ResNet18 to ResNet152), DenseNet [9] (DenseNet121 to DenseNet201), VGG [22] (VGG11 to VGG19, including with and without batch normalization), ViT [24, 6] and DeiT [27]. For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [37, 38], Celeba [15], Tiny ImageNet and ChestXRay [31, 16]. The information of these datasets can be find in Tab. 4. Specially, ChestXRay dataset has 14 labels and each of them is a binary classification task. We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay. For model ensemble, we consider ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer). Besides these two model ensemble techniques, whenever a model is considered to be merged into a greedy soup model, we ensemble it using either the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts). In order to exclude the impact of adopting the last layer from any single model, the “Ens Features” model uses the merging of the last fully connected (FC) layers from all merged models for classification.
@@
For all experiments in the paper, to keep fair comparisons, we adopt the official splits for each dataset. Also, we uniformly train 10 models and perform the model-merging and ensemble under exactly same conditions (learning rates, step strategies and optimizers, etc).
@@ 9 More Visualization of Model Weight Patterns
Figure 9: Average of all images within each class of Cifar100 dataset (all 100 classes).
Figure 10: Linear classifiers visualization for each class of Cifar100 dataset (all 100 classes).
Similar to the visualization in the main paper, we visualize the average of all images Fig. 9 and linear classifiers for all classes Fig. 10 within each class (all 100 classes) of Cifar100 dataset. We can easily and clearly observe the strong correlation between the average of images in a class and the learned weights.
@@
More visualization of the convolutional kernels on Cifar100 within deep learning models, e.g. Fig. 12 for ResNet50 and Fig. 13 for VGG19. For ResNet50 kernel visualization, we can observe differences or similarities between each learned patterns. For instance, Kernel 5 and kernel 9 are similar, which are responsible for detecting similar features and thus yield a higher measurement (e.g. inner product) between the two kernels; while Kernel 9 is very different with Kernel 45.The patterns inside ResNet50 is easier to be observed than VGG19, as ResNet50 has larger kernel size for the 1st layer.
Figure 14: Average of all images within each class of TinyImageNet dataset (the first 100 classes).
Figure 15: Linear classifiers visualization for each class of TinyImageNet dataset (the first 100 classes).
Figure 16: Average of all images within each class of TinyImageNet dataset (the last 100 classes).
Figure 17: Linear classifiers visualization for each class of TinyImageNet dataset (the last 100 classes).
@@
Figure 18: The 2 model merging/ensembling accuracy comparison between different architectures on Cifar100.
Figure 19: The 3 model merging/ensembling accuracy comparison between different architectures on Cifar100.
Figure 20: The 5 model merging/ensembling accuracy comparison between different architectures on Cifar100.
Figure 21: The 7 model merging/ensembling accuracy comparison between different architectures on Cifar100.
The 2 models, 3 models, 5 models and 7 models for model merging/ensembling accuracy comparison between different architectures on Cifar100 are shown via Fig. 18, Fig. 19, Fig. 20 and Fig. 21, respectively. They all show that the models with ensembling on logits generally receive the best results and uniform soups receive the worst results. Besides, model-merging performs much worse results for the ViT models and the gaps between uniform soups and ensemble on logits are enlarged with more models considered in.
Figure 22: Performance Gaps between Logits Ensemble and Greedy Soups Across Different Configurations (excluded the ViT models as the scales are too different) on Cifar100.
Fig. 22 shows the performance gaps between ensembling on logits with greedy soups. From it, we can observe similar phenomenon with the gaps between ensembling on logits with uniform soups, with more models merged, the larger gaps between ensemble on logits with greedy soups will be. But differently, the greedy soups break the going down trend with uniform soups as described in the main paper.
Table 5: Performance Comparison for VGG19 with 2 Models Across Different Datasets. For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
@@
Table 6: Performance Comparison for VGG19 with 3 Models Across Different Datasets. For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
@@
Table 7: Performance Comparison for VGG19 with 5 Models Across Different Datasets. For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
@@
Table 8: Performance Comparison for VGG19 with 7 Models Across Different Datasets. For the ChestXRay dataset, We adopt averged AUROC (Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation); while accuracy (0% to 100%) for the rest of datasets.
@@
Table 9: Comparison of FLOPs and Parameter Numbers for Various Models. The FLOPs are calculated based on 224 $\times$ 224 for general purpose, e.g. ImageNet. “ViT_Tiny” denotes for “ViT_Tiny_Patch16_224” and “DeiT_Tiny” stands for “DeiT_Tiny_Patch16_224”, respectively.
@@
We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in Tab. 10. We can also see the more models to be merged, the poorer results from uniform soups can get. Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed, it further demonstrates the model-merging between 2 or more models will not contribute to the performance.
@@ 11 More Experimental Results of Model Magnitude Changes
Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs. magnification factor on Cifar10 are shown via Fig. 27, Fig. 28, Fig. 29, Fig. 30 and Fig. 31, respectively. Moreover, we also visualize the uniform soup and logits ensemble accuracy vs. magnification factor across different model numbers on the Cifar10 dataset in Fig. 32. We can observe similar phenomenons. However, on the Cifar10, the gaps have different a peak point, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).